Ram moves in east direction at a speed of 6m/s and Shyam moves 30degree east of north at a speed of 6m/s. The magnitude of their reative velocity is?
 

Grade:11

2 Answers

Arun
5 years ago
Velocity of Ram = 6i
Velocity of Shyam = 6Sin30i + 6Cos30j

Relative velocity = 6i - 6Sin30i - 6Cos30j
= 6i - 3i - 3√3j
= 3i - 3√3j

Magnitude of relative velocity = sqrt(3^2 + (3√3)^2)
= 6 m/s
Aravind
5 years ago
Shyam moves 30east of North so the angle between them is 600,then
Vrs=√Vr^2+Vs^2-2 (Vr)(Vs)COS60
      =√6^2+6^2-2(6)(6)(1/2)
      =√36+36-36
      =√36
      =6m/s